OpenStudy (anonymous):

Find if the following conditions hold: r"(t) = a for all real t, r'(0) = 0, r(0) = b?

OpenStudy (anonymous):

r'(t)=integral of a=at+C but r'(0)=0 so C=0 so r'(t)=at r(t)=integral of at=at^2/2 +A but r(0)=b so A=b thus r(t)=at^2/2 + b

OpenStudy (anonymous):

Thanks! I see what to do now.